a) What is proposition?

b.) P and q are propositions given by p: f is an odd function and q: βˆ«π‘“(π‘₯)𝑑π‘₯=0π‘˜βˆ’π‘˜

verify whether or not π‘β†”π‘ž


1
Expert's answer
2021-02-24T07:33:42-0500
  1. A proposition, from a mathematical logic viewpoint, is a logical expression without free variables. From a viewpoint of "general" mathematics it is a statement meant to be proven.
  2. First let's look at "p \rightarrow q". Suppose that "f" is an odd function. Strictly speaking, a proposition q has a meaning only if "\int_{-k}^k f (x) dx" exists (i.e. f integrable), as we can have an odd function that is not Riemann-integrable/Lebesgue-integrable etc. Assuming that the expression "\int_{-k}^k f (x) dx" makes sense, the implication "p \rightarrow q" is true, as "\int_{-k}^k f(x)dx = \int_0^k f(x)dx+\int_{-k}^0 f(x)dx =0" by changing the variables in the second integral. The other direction "q \rightarrow p" depends on the class of "f" : if it is continuous, then it is true as "F(k) = \int_{-k}^k f(x)dx=0, F'(x) = f(x)+f(-x)=0" and thus "f(x) = -f(-x)", it is odd. If we make no hypothesis on "f", then it is obviously false (e.g. "f(0)=1, f(x)=0, x\neq 0") . Therefore for continuous functions this equivalence is true : "\forall f \in \mathcal{C}^0 (\mathbb{R}), \forall k\in \mathbb{R}, f(-x)=-f(x) \leftrightarrow \int_{-k}^k f=0".